given the endpoints of the diameter of the circle are (-8,4) and (-4,8). write the equation of the circle
Answer by lwsshak3(11628)

given the endpoints of the diameter of the circle are (-8,4) and (-4,8). write the equation of the circle
**
Standard form of equation for circle: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center, r=radius.
use midpoint formula to find center:
(x1+x2)/2, (y1+y2)/2
(-8-4)/2/(4+8)/2=(-6,6)
center: (-6,6)
use distance formula to find diameter:
d^2=(x1-x2)^2+(y1-y2)^2
d^2=(-8+4)^2+(4-8)^2
d^2=16+16=32
d=√32
radius=√32/2
r^2=32/4=8
Equation of given circle:
(x+6)^2+(y-6)^2=8